OpenStudy (anonymous):

Given that P = (5, 1) and Q = (13, 8), find the component form and magnitude of vector PQ

OpenStudy (mathmate):

Example: Given R(5,3), S(8,7), vector RS <8-5,7-3>=<3,4> ---- this is the component form magnitude RS =||RS|| = sqrt(3^2+4^2)=5 Direction of RS =arctan(4,3) (but watch quadrant, most arctan function give solution in -pi/2 to pi/2 only) = 53.13 deg. approx.